# Leveraging Large Language Models to Address Data Scarcity in Machine Learning: Applications in Graphene Synthesis

### v1

@@ 3. K-means Binning: Groups similar values into $k$ clusters
The discretization techniques are validated using the SVM classifier with the same parameters used for the featurization testing. To keep track of the encoded labels used in the discretization step, we create a bin map that assigns each bin with it’s respective range and can be retrieved later to uncover the meaning of those labels. Supplementary Figure 8 shows a data frame with discretized variable space and it’s corresponding bin map.
@@
The effect of the number of bins on the accuracy of equal-width and equal-frequency binning techniques is shown in Figure Supplementary Figure 9.
@@
Supplementary Figure 10: Function for K-means binning with the selection of optimal number of bins.
@@
The choice of the ML architecture is primarily related to the learning capacity of the model. In the main manuscript, we present the learning rates of different ML models and conclude that the support vector machine (SVM) classifier is the optimal choice for the current dataset. For the binary classification problem, our dataset contains 83 instances of monolayer and 81 instances of multilayer graphene. For the ternary classification, the multilayer category can be further divided into 31 bilayer and 50 multilayer instances. The number of layers is label-encoded as follows: 0 for monolayer, 1 for bilayer, and 2 for multilayer (for ternary classification). The data are split into an 80-20 train-test ratio, with extensions to a 90-10 split for incremental learning curve analysis.
@@
We use the learning_curve function from the scikit-learn library with training sizes obtained via np.linspace (0.1,1,5) with 5-fold cross-validation (cv=5) to compute the learning curve. This technique involves splitting the dataset into five subsets of training sizes and five folds of validation sets, where the model is trained on each training subset and tested on each corresponding validation set. The training scores are obtained by evaluating the model on different training data sets for each fold. The test score is obtained by evaluating the model on the respective validation fold. Averaging the training and test scores, we get the learning and cross-validation curves, respectively. A clear description of this method can be found in the scikit-learn library or the Supplementary file Python notebook. The binary SVM classifier learning curves for different prompt-based predictions using the SE-4 featurization technique are shown in Supplementary Figure 11. Additionally, the same techniques are applied to the datasets imputed by the MAP and CITE prompts and their learning curves are obtained using the SVM classifier as shown in Supplementary Figure 12.
@@
where $\alpha_{i}$ is the Lagrangian multiplier for the $i^{\text{th}}$ data point, and $K(\mathbf{x}_{i},\mathbf{x}_{j})$ is the kernel function that maps input data into a higher-dimensional space.The regularization parameter (C) plays a crucial role in adjusting the trade-off between maximizing the margin and minimizing classification errors, particularly in cases of scarce data with biased attributes. While custom regularization techniques are beyond the scope of this study, we recommend exploring such approaches as a potential strategy to enhance the accuracy of an SVM classifier.
@@
We optimize the SVM hyperparameters by exhaustively searching over a predefined parameter grid for the regularization parameter $C$ (0.1, 1, 10, 100), the kernel coefficient $\gamma$ (‘‘scale’’, ‘‘auto’’, 0.01, 0.1, 1, 10), and the kernel type (RBF, linear, polynomial). This search is carried out using GridSearchCV with 5-fold cross-validation, where each hyperparameter combination is evaluated by its average accuracy across the folds. The best-performing set of parameters is then automatically selected, and the final SVM model is retrained on the entire training set before being tested on the held-out data.
@@
OpenAI specifies the requirement of a JSONL format dataset consisting of a System, User, and Assistant prompt with the first one requiring general instructions, the second one corresponding to the input prompt, and the final one being the output or response desired from the model. We write a simple function that specifies general classifier instruction for CVD and the name of the attributes, the user prompt consisting of the actual values of these attributes, and the assistant prompt specifying integers- 0,1, or 2- corresponding to monolayer, bilayer, or multilayer graphene, respectively. We divide the dataset into training and validation split (80-20) beforehand and specify them during the job creation process. The datasets and data preparation scripts are provided in our Github page.
@@
We submit our training and validation datasets to the OpenAI job module, along with the specification of (a) the number of epochs to 3, (b) the learning rate multiplier to 1.0 (c) the batch-size to 1, and (d) the model suffix (e.g., 4o or 4o-mini). The fine-tuning job terminates upon full validation over the three epochs, and the results are obtained as outputs of the job.
@@
The token prediction accuracy does not represent the graphene layer classification accuracy but only the accuracy with which the output embeddings are predicted by ChatGPT. We write a function to get the classification results on the validation set. The classification results are then used to calculate performance metrics including the accuracy, precision, recall, and F1 score. Supplementary Table 5 shows the results of different fine-tuned models with the cross-entropy accuracy as well as the actual classification accuracies.
@@
Supplementary Table 5: Comparison of cross-entropy accuracy for token prediction versus graphene layer classification accuracy, precision, recall and F1 score.

### v2

@@ 3. K-means Binning: Groups similar values into $k$ clusters
The discretization techniques are validated using the SVM classifier with the same parameters used for the featurization testing. To keep track of the encoded labels used in the discretization step, we create a bin map that assigns each bin with its respective range, which can be retrieved later to uncover the meaning of those labels. Supplementary Figure 8 shows a data frame with discretized variable space and its corresponding bin map.
@@
The effect of the number of bins on the accuracy of equal-width and equal-frequency binning techniques is shown in Supplementary Figure 9.
@@
Supplementary Figure 10: Function for K-means binning with the selection of an optimal number of bins.
@@
The choice of the ML architecture is primarily related to the learning capacity of the model. In the main manuscript, we present the learning rates of different ML models and conclude that the support vector machine (SVM) classifier is the optimal choice for the current dataset. For the binary classification problem, our dataset contains 83 monolayer and 81 instances of multilayer graphene. The multilayer category can be further divided into 31 bilayer and 50 multilayer instances for the ternary classification. The number of layers is label-encoded as follows: 0 for monolayer, 1 for bilayer, and 2 for multilayer (for ternary classification). The data are split into an 80-20 train-test ratio, with extensions to a 90-10 split for incremental learning curve analysis.
@@
We use the learning_curve function from the scikit-learn library with training sizes obtained via np.linspace (0.1,1,5) with 5-fold cross-validation (cv=5) to compute the learning curve. This technique involves splitting the dataset into five subsets of training sizes and five folds of validation sets, where the model is trained on each training subset and tested on each corresponding validation set. The training scores are obtained by evaluating the model on different training data sets for each fold. The test score is obtained by evaluating the model on the respective validation fold. Averaging the training and test scores, we get the learning and cross-validation curves, respectively. A clear description of this method can be found in the scikit-learn library or the Supplementary file Python notebook. The binary SVM classifier learning curves for different prompt-based predictions using the SE-4 featurization technique are shown in Supplementary Figure 11. Additionally, the same techniques are applied to the datasets imputed by the MAP and CITE prompts, and their learning curves are obtained using the SVM classifier as shown in Supplementary Figure 12.
@@
where $\alpha_{i}$ is the Lagrangian multiplier for the $i^{\text{th}}$ data point, and $K(\mathbf{x}_{i},\mathbf{x}_{j})$ is the kernel function that maps input data into a higher-dimensional space.The regularization parameter (C) plays a crucial role in adjusting the trade-off between maximizing the margin and minimizing classification errors, particularly in scarce data with biased attributes. While custom regularization techniques are beyond the scope of this study, we recommend exploring such approaches as a potential strategy to enhance the accuracy of an SVM classifier.
@@
We optimize the SVM hyperparameters by exhaustively searching over a predefined parameter grid for the regularization parameter $C$ (0.1, 1, 10, 100), the kernel coefficient $\gamma$ (‘‘scale’’, ‘‘auto’’, 0.01, 0.1, 1, 10), and the kernel type (RBF, linear, polynomial). This search uses GridSearchCV with 5-fold cross-validation, where each hyperparameter combination is evaluated by its average accuracy across the folds. The best-performing set of parameters is then automatically selected, and the final SVM model is retrained on the entire training set before being tested on the held-out data.
@@
OpenAI specifies the requirement of a JSONL format dataset consisting of a System, User, and Assistant prompt, with the first one requiring general instructions, the second one corresponding to the input prompt, and the final one being the output or response desired from the model. We write a simple function that specifies general classifier instruction for CVD and the name of the attributes, the user prompt consisting of the actual values of these attributes, and the assistant prompt specifying integers- 0,1, or 2- corresponding to monolayer, bilayer, or multilayer graphene, respectively. We divide the dataset into training and validation splits (80-20) beforehand and specify them during the job creation. The datasets and data preparation scripts are provided on our GitHub page.
@@
We submit our training and validation datasets to the OpenAI job module, along with the specification of (a) the number of epochs to 3, (b) the learning rate multiplier to 1.0, (c) the batch size to 1, and (d) the model suffix (e.g., 4o or 4o-mini). The fine-tuning job terminates upon full validation over the three epochs, and the results are obtained as outputs of the job.
@@
The token prediction accuracy does not represent the graphene layer classification accuracy but only the accuracy with which the output embeddings are predicted by ChatGPT. We write a function to get the classification results on the validation set. The classification results are then used to calculate performance metrics, including the accuracy, precision, recall, and F1 score. Supplementary Table 5 shows the results of different fine-tuned models with the cross-entropy accuracy and the actual classification accuracies.
@@
Supplementary Table 5: Comparison of cross-entropy accuracy for token prediction versus graphene layer classification accuracy, precision, recall, and F1 score.